what work needs to be shown in hte problem stating, "determine the value of k for which the system of linear equations has no solution."
4x + 5y=12
8x + ky=13

Answer by ankor@dixie-net.com(22740)

"determine the value of k for which the system of linear equations has no solution."
4x + 5y=12
8x + ky=13
:
Multiply the 1st equation by 2
8x + 10y = 24
8x + ky = 13
:
Clearly if k = 10, there is no solution